A conducting cylinder of radius 3.25 cm and length 5.6 cm has a total charge of 5.5 x 10-9 C distributed uniformly on its surface area. Find for the potential at its surface.


1
Expert's answer
2022-02-17T11:37:14-0500

Answer

conducting cylinder of radius

R=3.25 cm and

length L=5.6 cm

total charge of q=5.5 x 10-9 C


Potential is given by

"V=\frac{\lambda}{2\pi \epsilon_0}\ln(\frac{r}{R})"

For surface r=R

So logrithm term become zero. So potential at surface will be zero.